OpenStudy (anonymous):

I need help on how to simplify complex rational expressions & linear equations

OpenStudy (anonymous):

This is the complex rational expression I'm working with,

OpenStudy (anonymous):

And this is the linear equation I'm working with,

OpenStudy (tkhunny):

1) Reciprocal and Multiply. 2) Before actually multiplying, factor everything and see if your life is improved.

OpenStudy (anonymous):

So multiply 6 by 5 / x^2 - 36?

OpenStudy (tkhunny):

6? \(\dfrac{\dfrac{1}{x+6}}{\dfrac{5}{x^{2}-36}} = \dfrac{1}{x+6}\cdot\dfrac{x^{2}-36}{5}\) Now, it is time to factor where you can.

OpenStudy (anonymous):

Ah, when you said reciprocal I was thinking you meant 6 / 1

OpenStudy (tkhunny):

:-) This is why we talk! How's that factoring going?

OpenStudy (anonymous):

Hmm, I'm not sure if I'm factoring properly, like this.?